OpenStudy (anonymous):

For f(x)=5+1, find f(-4)

OpenStudy (anonymous):

I think you mean f(x)=5x+1. That would be f(-4)=5(-4)+1. f(-4)=-20+1 f(-4)=-19

OpenStudy (anonymous):

If you typed the whole question of f(x)=5+1, then f(x)=6 y=6 If x=-4 and y=6